OpenStudy (woahning):

the midpoint of uv is (5 -11). the coordinates of one endpoint are u(3,5). Find the coordinates of endpoint v, Can someone please help and show the work because I just don't understand this at all

TheSmartOne (thesmartone):

\[midpoint=(\frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1} +y _{2}}{2})\]

TheSmartOne (thesmartone):

So\[\frac{ 3+x }{ 2}=5\] \[3+x=10\] x=7 and \[\frac{ 5+y }{ 2 }=-11\] \[5+y=-22\] \[y=-27 \] So the points are \[(7,-27)\]